akecaption[2]

mtempboxa1: 2mtempboxa ¿1: 2

(revised 1/29/99)

Compton Scattering

Advanced Laboratory, Physics 407 University of Wisconsin Madison, Wisconsin 53706

\underline{Abstract}

The Compton scattering of the 662 keV gamma rays from the decay of Cs^{137} is measured using a Sodium Iodide detector. The scattered energy and the differential cross section are both measured as a function of scattering angle, and the results are compared to the full relativistic quantum theory of radiation.